Two students are using a measuring tape to find the length of a room.The tape reads 13 ft 6.25 inches.However, the student at one end is holding the tape 9 in. higher than the student at the other end. What is the difference between their measurement and the true length?
Answer by scott8148(6628)

the tape is the hypotenuse of a right triangle , while the 9 in and the true length are the two sides

(13 ft 6.25 in)^2 = (t)^2 + (9 in)^2

162.25^2 = t^2 + 9^2 ___ t^2 = 162.25^2 - 9^2 ___ t = sqrt(162.25^2 - 81)